OpenStudy (anonymous):

3 kingkos Group Neophyte given that f(x) = 2x + 1 and g(x) = x^2 -2.find (f o g) (x + 2). choices a. 2x^2 - 8x + 5 c. 2x^2 + 8x - 5 b. 2x^2 + 8x +5 d.2x^2 - 8x - 5 pls help for my H.w

OpenStudy (anonymous):

\[g(x+2)=(x+2)^2-2=x^2+4x+4-2=x^2+4x+2\]

OpenStudy (anonymous):

what letter is the correct answer all are the same just a sign of positive and negative

OpenStudy (anonymous):

\[f(g(x+2))==2(x^2+4x+2)+1\] \[=2x^2+8x+4+1\] \[=2x^2+8x+5\]

OpenStudy (anonymous):

yes! thank you! last 1 sir

OpenStudy (anonymous):

actually i think it is b yes?

OpenStudy (anonymous):

correct sir thank you

OpenStudy (anonymous):

sir i have another question